I want the answer in typing if the answer in hand writing I will give dislike

In typing no handwriting

IN TYPING PLEAEEEEEEEEEEEEEEEEEESE !

Discussion Question: in typing -

Discuss a scenario (with examples & Chemical Equations/Calculations) which involves the application of the two gas laws that are given below.

1- Boyles law &

2- Charles law
Answer #1

1) Boyles Law

Definition: Boyles law states that the pressure exerted by a gas(of a given mass, kept at a constant temperature) is inversely proportional to the volume occupied by it

Equation is

P = k* (1 /V) ----> PV = k

Examples of Boyles Law

1) If a scuba driver rapidly ascends from a deep zone towards the surface of water, the decrease in the pressure can cause the gas molecules in his/ her body to expand. These gas bubbles can go on to cause damage to the driver's organs and can also result in death. This expansion of gas caused by the asscension of the scuba driver is an example of Boyles Law.

2) Deep sea fish that die after reaching the surface of the water ( due to the expansion of dissolved gases in their blood)

Calculation

A fixed amount of gas occupies a volume of 1 L and exerts a pressure of 400kpa on the walls of its container. What would be the pressure exerted by the gas if it is completely transferred in to new container having a volume of 3 litres.

Given

Initial Volume (V1) = 1L

Initial pressure (P1) = 400kPa

Final volume (V2) = 3L

As per Boyles Law, P1V1 = P2V2 ----> P2 =( P1V1) /V2

P2 = (1L* 400KPa) / 3L

=133.3KPa

Therefore the gas exerts the pressure of 133.3 kPa on the walls of the 3 litre container.

2) Charles Law

Definition: It states that when pressure is held constant , the volume of a fixed amount of dry gas is directly proportional to its absolute temperature.

Equation is

V1/T1 = V2/T2

V1 and V2 are the initial volume and final volume respectively and T1 and T2 are the initial and final temperatures respectively

Examples of Chrles Law

1) In cold weather or environment, balls and helium balloons shrink.

2) In bright sunlight, the inner tubes swell up.

Calculation:

A gas occupies a volume of 400cm3 at 0 degrre celsius 780mm Hg. What volume (in litres) will it occupy at 80 degree celsius and 780mm Hg.

Given

V1 = 400cm3

V2 =?

T1 = 0 degree ceisius = 0 + 273 = 273 K

T2 = 80 degree celsius = 80 + 273 = 353 K

Here the pressure is constant and only the temperature is changed.

Using Charles Law,

(V1) /(T1) = (V2) / (T2)

400 / 273 = V2 / 353

V2 = 400 X 353 / 273

V2 = 517.21cm3

1 cubic centimeter = 0.001 litre = 1 X 10-3 litre

517.21 cubic centimeter = 517.21 X 10-3 = 0.517litres.